40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1963     Entry time: Mon Oct 30 16:15:09 2017
Author: Craig, awade 
Type: DailyProgress 
Category: BEAT 
Subject: PLL locked at ~82 MHz 

awade and I managed to lock both cavities today after diagnosing the acromag communications issues.  Some stats:

South Slow Voltage: -6.12 V

North Slow Voltage: 3.46 V

Vaccan Temp: 30 Celcius

Marconi FM Devn: 25 kHz

Marconi RF Lvl: +12.6 dBm

Trans Beatnote Frequency: 82.4 MHz (and drifting up at ~5 kHz/s)

Channels dumped into acromag1 at /home/controls/CTNWS/data/20171030_161251_ChannelDump.txt

awade is currently playing with the Marconi to try and get the FM Devn down to around 1 kHz so we can get a spectrum that is not limited by actuation noise.  Beatnote frequency drift is settling down significantly as I type this, so it could be possible soon.


∆f/f = ∆L/L for the Fabry Perot cavity.  f = 282 THz, L = 3.68 cm, and ∆f = ~ 5 kHz/s implies that ∆L ~ 6.5e-13 m/s for our relative cavity length changes. 

From elog 1874, we have ∆T ~ ∆L / (alpha * mirror thickness), where alpha = 5.5e-7 [1/Kelvin] and mirror thickness = 6.35 mm.  We get a relative temperature drift of ∆T ~ 0.2 mK / s

We seem to have underestimated how stable our cavities must be to temperature drift.  In any case our beatnote is still drifting rapidly despite the vaccan stabilization.

ELOG V3.1.3-